OpenStudy (dls):

The pH of a monobasic acid HA is 4.50.Calculate concentration of everything and Kc.

OpenStudy (dls):

I got pH=3.3*10^-5 and pOH=3.3*10^-10 hopefully..just an attempt..

OpenStudy (anonymous):

what do you mean about Kc!!!!!!!!!!!!!!

OpenStudy (aaronq):

use the henderson hasselbach equation for the Kc... although i think you may be missing information.

OpenStudy (jfraser):

you might be missing the overall concentration of acid, but knowing pH gets you pOH, [H+], [OH-]. You still need Ka, and Kb